# Rate of epidemic spreading on complex networks

### v3

@@ Abstract
The initial phase of an epidemic is often characterized by an exponential increase in the number of infected individuals. In this Letter, we predict the exponential spreading rate of an epidemic on a complex network. We first find an expression of the reproduction number for a network, based on the degree distribution, the network assortativity, and the level of clustering. We then connect this reproduction number and the disease infectiousness to the spreading rate. Our result holds for a broad range of networks, apart from networks with very broad degree distribution, where no clear exponential regime is present. Our theory bridges the gap between classic epidemiology and the theory of complex network, with broad implications for model inference and policy making.
@@
Epidemic modeling permits rationalization of evidences and provides crucial information for policy making. Epidemiological models should incorporate the key features that determine epidemic spreading. For example, social structures play a crucial role in epidemic spreading Buckee et al. (2021), and in the effectiveness of intervention strategies Patwardhan et al. (2023); Granell and Mucha (2018). In these social structures, the number of contacts per individual is highly variable. For instance, individuals with most contacts contributed disproportionately to the spreading of COVID-19 Kumar et al. (2020); Sneppen et al. (2020).
@@
Social structures can be conveniently modeled by means of complex networks Newman (2018), in which nodes represent individuals and links denote pairs of individuals that are in contact. A success of network theory has been to determine the epidemic threshold on networks Pastor-Satorras and Vespignani (2001); Newman (2002); Pastor-Satorras et al. (2015), i.e., the transition point from a localized outbreak to a widespread epidemic. This theory has shown that scale-free networks are extremely vulnerable to epidemics Pastor-Satorras and Vespignani (2001); Barthélemy et al. (2005); Pastor-Satorras et al. (2015); Pósfai and Barabási (2016) . Clustering, i.e., the tendency of nodes to form tightly connected communities, also affects epidemic spreading. Many real-world networks possess both heavy-tailed degree distribution and high level of clustering Ravasz and Barabási (2003).
@@
Another key feature of an epidemic is its infectiousness, i.e., the rate at which an infected individual infects its contacts. The infectiousness incorporates multiple factors, such as the viral load, social behavior, and possibly environmental factors Grassly and Fraser (2008). Because of changes in these factors, the infectiousness of an individual changes over time in a way that is particular to each disease. This time-dependence profoundly impacts how a disease spreads.
@@
Social structures and individual infectiousness are crucial in determining the initial stage of an epidemic, in which most individuals are susceptible and the number of infected individuals usually grows exponentially. This stage is of central importance for novel epidemics, or during emergence of novel variants, as it often requires rapid policy making. At later stages, other factors become relevant, such as individual recovery time, their behavioral response Weitz et al. (2020), and the possible occurrence of reinfections.
@@
In this Letter, we theoretically predict the exponential rate of epidemic spreading on a complex network. Our theory quantifies the impact of main properties of the network, such as assortativity and clustering, on the reproduction number of the epidemic. We further link the reproduction number and the individual infectiousness with the exponential spreading rate. Our approach leads to precise predictions for a broad range of networks.
Epidemic spreading in well-mixed populations. We begin by summarizing how the exponential spreading rate of an epidemic is determined in well-mixed populations Grassly and Fraser (2008). In this case, the cumulative number $I(t)$ of infected individuals at time $t$ is governed by the renewal equation
@@
where the basic reproduction number $R_{0}$ represents the average number of secondary infections by each infected individual, and the dot denotes a time derivative. The function $w(\tau)$ represents the probability density that a given secondary infection occurs after a time $\tau$ from the primary one. Equation (1) is linear because of the assumption that the epidemic is in the exponential stage and thus far from saturation. We now substitute the ansatz $I(t)\sim e^{\Lambda t}$ into Eq. (1) and take the limit of large $t$, obtaining the Euler-Lotka equation
@@
Over the course of the epidemic, each individual can be infected at most once and never formally recovers, although their infectiousness may wane over time. Our model can thus be see as a non-Markovian Susceptible-Infected model. The time dependence of $\lambda(\tau)$ is important to model real-world epidemics. For example, the infectiousness of COVID-19 markedly increases for 3–6 days and then decays for about 11 days Hakki et al. (2022), and is therefore poorly described by a Markovian model with constant $\lambda(\tau)$.
@@
In the model, the number of infected individuals grows in time until it saturates (Fig. 1c). An alternative to the temporal perspective is to represent the epidemic in generations (Fig. 1d). We call $Z_{n}$, $n=0,1,\dots$ the average number of infected individuals in the $n$th generation. The initial generation, $n=0$, is constituted by the patient zero, $Z_{0}=1$. In this representation, the infected individuals form a tree (Fig. 1e).
@@
Figure 1: (a) Epidemic spreading on a network. Individuals (nodes) can be infected (red) or healthy (blue). Infected individuals spread the infection at a rate $\lambda(\tau)$ to each of their neighbors, where $\tau$ is the time since they were infected. If neighbors have already been infected, transmission does not result in an infection. (b) Time-dependent spreading rate $\lambda(\tau)$ and resulting distribution $\psi(\tau)$ of infection times. (c) An epidemic trajectory. The exponential phase ends in a saturation phase, where a sizeable fraction of the individuals are infected. (d) Epidemic tree. The epidemic initiates with one infected individual (patient zero). Subsequent infections are represented by continuous lines. Dashed lines represents generations.
Reproduction number on a network. The basic reproduction number $R_{0}$ for a network is the average number of links of an infected individual in the early stage of the epidemic, excluding the one from whom the infection came from. The effective reproduction number $R$ is the average number of infections caused by an infected individual. In contrast with $R_{0}$, $R$ does not include contacts that are infected by other individuals. This difference is crucial in clustered networks.
@@
whose elements are the average numbers of susceptible individuals of degree $k$ connected to a node of degree $k^{\prime}$. In Eq. (3), $P(k|k^{\prime})$ is the probability that a neighbor of a node of degree $k^{\prime}$ has degree $k$ and $m_{kk^{\prime}}$ is the average number of triangles a $kk^{\prime}$-link is part of. The term proportional to $m_{kk^{\prime}}$ discounts for contacts of a given individual that were already infected by someone else, due to the presence of triangles. In Eq. (3), we neglect higher order loops such as squares. The reproduction matrix has been previously used to study percolation on networks Serrano and Boguná (2006).
@@
At the first generation, we obtain $Z_{1}=\bar{k}$. For large $n$, we have $Z_{n}\sim R^{n}$, where we identify the reproduction number $R$ as the leading eigenvalue of $M_{kk^{\prime}}$. The associated right eigenvector $v_{k^{\prime}}$ represents the degree distribution of infected individuals sup .
@@
where the bar represents an average over the degree distribution. Equation (5) is a classic result Pastor-Satorras and Vespignani (2001); Callaway et al. (2000); Newman (2002). We seek to extend it to the general case of clustered and correlated networks. We quantify degree correlations between connected nodes by the assortativity coefficient $r=1/\sigma_{r}^{2}(\sum_{kk’}kk^{\prime}P(k,k’)-(\overline{k^{2}}/\overline{k})^{2})$,with $\sigma_{r}^{2}=\overline{k^{3}}/\overline{k}-(\overline{k^{2}}/\overline{k})^{2}$ sup . We find that, for small clustering and assortativity coefficients, the effective reproduction number is expressed by
$R_{\mathrm{pert}}=(1-r)(R_{0}-m_{1})+r\left(\frac{\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}}{R_{0}-m_{1}}\right)\,,$ (6)
where $c_{k}$ is the probability that two neighbors of a node of degree $k$ are themselves connected, $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$ is the average number of triangles a link is part of, and $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$. We obtained Eq. (6) by a perturbative expansion in $r$ sup .
We compare $R_{\mathrm{pert}}$ given by Eq. (6) with the maximum eigenvalue $R$ of the matrix $M_{kk^{\prime}}$ for different network models, see Table 1. The relative difference between $R$ and $R_{\mathrm{pert}}$ is within $2\%$ for all the models we considered, except for Barabási-Albert with parameter $m=1$ ($20.6\%$) and $m=2$ ($7.1\%$), where $m$ represents the number of nodes a new node attaches to. Approximating $R$ with $R_{0}$ as given by Eq. (5) leads to substantially larger errors, supporting that clustering and assortativity play an important role. Here and in the following, when computing $R_{\mathrm{pert}}$ using Eq. (6) and $R_{0}$ using Eq. (5) for a given network, we always interpret the averages as empirical averages. With this choice, moments of the degree distribution remain finite even in very heterogeneous networks. In practical cases, Eq. (6) predicts that $R_{\mathrm{pert}}$ increases with the assortativity $r$. Accordingly, assortativity lowers the epidemic threshold, as previously known for the unclustered case Goltsev et al. (2012). However, this dependence is usually weak, unless the degree distribution has heavy tails. In contrast, clustering substantially reduces $R_{\mathrm{pert}}$. For this reason, we find that $R_{0}$ overestimates $R$ in all the synthetic networks we considered.
Table 1: Reproduction numbers. Here, $R$ is the numerically computed leading eigenvalue of Eq. (3). The expressions of $R_{0}$ and $R_{\mathrm{pert}}$ are given by Eq. (5) and Eq. (6), respectively. To assess the impact of assortativity and clustering, we also compute $R_{\mathrm{pert}}$ for $r=0$ and for $\bar{c}=\sum_{k}c_{k}P(k)=0$. BAm: Barabási-Albert with parameter $m$. $\text{ER}_{d}$: Erdős-Rényi with mean degree $d$. WSk,p: Watts-Strogatz starting from a 1D lattice with $k$ nearest neighbors and rewiring probability $p$.
@@
Network Euler-Lotka equation. We now use our estimate of $R$ to predict how epidemics spread in time. The average total number ${I}_{k}(t)$ of infected individual of degree $k$ at time $t$ is governed by the renewal equation

### v4

# Exponential rate of epidemic spreading on complex networks

@@ Abstract
The initial phase of an epidemic is often characterized by an exponential increase in the number of infected individuals. In this paper, we predict the exponential spreading rate of an epidemic on a complex network. We first find an expression of the reproduction number for a network, based on the degree distribution, the network assortativity, and the level of clustering. We then connect this reproduction number and the disease infectiousness to the spreading rate. Our result holds for a broad range of networks, apart from networks with very broad degree distribution, where no clear exponential regime is present. Our theory bridges the gap between classic epidemiology and the theory of complex networks, with broad implications for model inference and policy making.
I Introduction
@@
Epidemic modeling permits the rationalization of evidence and provides crucial information for policy making. Epidemiological models should incorporate the key features that determine epidemic spreading. For example, social structures play a crucial role in epidemic spreading Buckee et al. (2021) and in the effectiveness of intervention strategies Patwardhan et al. (2023); Granell and Mucha (2018). In these social structures, the number of contacts per individual is highly variable. For instance, individuals with most contacts contributed disproportionately to the spreading of COVID-19 Kumar et al. (2020); Sneppen et al. (2020).
@@
Social structures can be conveniently modeled by means of complex networks Newman (2018), in which nodes represent individuals and links denote pairs of individuals that are in contact. A success of network theory has been to determine the epidemic threshold on networks Pastor-Satorras and Vespignani (2001); Newman (2002a); Pastor-Satorras et al. (2015), i.e., the transition point from a localized outbreak to a widespread epidemic. This theory has shown that scale-free networks are extremely vulnerable to epidemics Pastor-Satorras and Vespignani (2001); Barthélemy et al. (2005); Pastor-Satorras et al. (2015); Pósfai and Barabási (2016). Clustering, i.e., the tendency of nodes to form tightly connected communities, also affects epidemic spreading. Many real-world networks possess both heavy-tailed degree distribution and high levels of clustering Ravasz and Barabási (2003).
@@
Another key feature of an epidemic is its infectiousness, i.e., the rate at which an infected individual infects its contacts. The infectiousness incorporates multiple factors, such as viral load, social behavior, and possibly environmental factors Grassly and Fraser (2008). Because of changes in these factors, the infectiousness of an individual changes over time in a way that is particular to each disease. This time dependence profoundly impacts how a disease spreads.
@@
Social structures and individual infectiousness are crucial in determining the initial stage of an epidemic, in which most individuals are susceptible and the number of infected individuals usually grows exponentially. This stage is of central importance for novel epidemics, or during the emergence of novel variants, as it often requires rapid policy making. At later stages, other factors become relevant, such as individual recovery time, their behavioral response Weitz et al. (2020), and the possible occurrence of reinfections.
@@
In this article, we theoretically predict the exponential rate of epidemic spreading on a complex network. Our theory quantifies the impact of the main properties of the network on the reproduction number of the epidemic and then links this reproduction number with the exponential spreading rate. We shall see that our approach leads to precise predictions for a broad range of networks.
The manuscript is organized as follows. Section II reviews classic concepts in epidemiology, in particular the definition of the reproduction number for well-mixed populations and its relation with the exponential spreading rate. Section III introduces our model of epidemic spreading on networks. Section IV discusses definitions of the reproduction number for networks, and shows how the reproduction number can be estimated. Section V contains our main result: the network Euler-Lotka equation, linking the reproduction number with the exponential spreading rate. Section VI extends our results to weighted networks. Section VII contains a discussion and future perspectives .
II Epidemic spreading in well-mixed populations
We begin by summarizing how the exponential spreading rate of an epidemic is determined in well-mixed populations Grassly and Fraser (2008). We call $I(t)$ the average cumulative number of infected individuals at time $t$. The average number of newly infected individuals $\dot{I}(t)$ is governed by the renewal equation
@@ II Epidemic spreading in well-mixed populations
where the basic reproduction number $R_{0}$ represents the average number of secondary infections by each infected individual. The function $\omega(\tau)$ represents the probability density that a given secondary infection occurs after a time $\tau$ from the primary one. Equation (1) is linear because of the assumption that the epidemic is in the exponential stage and thus far from saturation. We now substitute the ansatz $I(t)\sim e^{\Lambda t}$ into Eq. (1) and take the limit of large $t$, obtaining the Euler-Lotka equation
@@ III Model
We assume that, over the timescale of interest, each individual can be infected at most once and never formally recovers, although their infectiousness may wane over time. Our model can thus be seen as a non-Markovian susceptible-infected (SI) model. We however remark that, in this non-Markovian framework, the distinction between SI and a susceptible-infected-recover (SIR) is only formal, as individuals whose infectivity $\lambda(\tau)$ has decayed to zero can be practically considered as recovered.
@@
The time dependence of $\lambda(\tau)$ is important to model real-world epidemics. For example, the infectiousness of COVID-19 markedly increases for three to six days and then decays for about 11 days Hakki et al. (2022), and is therefore poorly described by a Markovian model with constant $\lambda(\tau)$.
@@
In the model, the number of infected individuals grows in time until it saturates [Fig. 1(c)]. An alternative to the temporal perspective is to represent the epidemic in generations [Fig. 1(d)]. We call $Z_{n}$, $n=0,1,\dots$ the average number of infected individuals in the $n$th generation. The initial generation, $n=0$, is constituted by the patient zero, $Z_{0}=1$. In this representation, the infected individuals form a tree [Fig. 1(d)].
@@
Figure 1: (a) Epidemic spreading on a network. Individuals (nodes) can be infected (red) or healthy (blue). Infected individuals spread the infection at a rate $\lambda(\tau)$ to each of their neighbors, where $\tau$ is the time since they were infected. If neighbors have already been infected, transmission does not result in an infection. (b) Time-dependent spreading rate $\lambda(\tau)$ and resulting distribution $\psi(\tau)$ of infection times. (c) An epidemic trajectory. The exponential phase ends in a saturation phase, where a sizable fraction of the individuals are infected. (d) Epidemic tree. The epidemic initiates with one infected individual (patient zero). Subsequent infections are represented by continuous lines. Dashed lines represent generations.
IV Reproduction number on a network
IV.1 Case $p_{\psi}=1$
In this section, we discuss the definition of the reproduction number for a network. We first focus, for simplicity, on the case $p_{\psi}=1$, where individuals infect all of their contacts and the epidemic eventually spreads through the network, if the network is connected.
The definition of the basic reproduction number $R_{0}$ for networks is slightly different from that in well-mixed populations. In well-mixed populations, $R_{0}$ is defined as the average number of secondary infections produced by a single infected individual in a fully susceptible population Grassly and Fraser (2008). Extending this definition to networks requires a specification of how the infecting node is selected. A node chosen uniformly at random will on average produce $\bar{k}=\sum_{k}kP(k)$ secondary infections, where $P(k)$ is the probability that the selected node has degree $k$. However, on a network infecting nodes are not chosen uniformly at random, but are rather infected by one of their neighbors. Consequently, susceptible nodes with higher degree are more likely to be infected. We therefore define $R_{0}$ as the average number of links of an infected individual in the early stage of the epidemic, excluding the link from which the considered individual received the infection. An equivalent definition is that $R_{0}$ is the average degree (minus one) of the neighbor of a random node Newman (2002a).
We also define the effective reproduction number $R$ as the average number of infections caused by an infected individual in the early stage of epidemics. In contrast to $R_{0}$, $R$ does not include contacts that are infected by other individuals. This distinction is necessary for networks, at variance with well-mixed populations. The reason is that, in well-mixed populations, most contacts of an individual are susceptible if the fraction of infected individuals is low. However, this is not necessarily the case in networks, as the neighborhood of an infected individual can be locally saturated even if the global fraction of infected individuals is low.
@@ IV.1 Case $p_{\psi}=1$
whose elements are the average numbers of susceptible individuals of degree $k$ connected to a node of degree $k^{\prime}$. In Eq. (3), $P(k|k^{\prime})$ is the probability that a neighbor of a node of degree $k^{\prime}$ has degree $k$ and $m_{kk^{\prime}}$ is the average number of triangles a $kk^{\prime}$-link is part of. The term proportional to $m_{kk^{\prime}}$ discounts for contacts of a given individual that were already infected by someone else, due to the presence of triangles. In Eq. (3), we neglect higher-order loops such as squares. The reproduction matrix has been previously used to study percolation on networks Serrano and Boguná (2006).
@@
At the first generation, we obtain $Z_{1}=\bar{k}$. For large $n$, we have $Z_{n}\sim R^{n}$, where we identify the reproduction number $R$ as the leading eigenvalue of $M_{kk^{\prime}}$. The associated right eigenvector $v_{k^{\prime}}$ represents the degree distribution of infected individuals, see Appendix A.
@@
where the bar represents an average over the degree distribution. Equation (5) is a classic result Pastor-Satorras and Vespignani (2001); Callaway et al. (2000); Newman (2002a).
For clustered and correlated networks, $R_{0}$ is still given by the right-hand side of Eq. (5), while the value of $R$ may in general differ. To estimate this value, we quantify degree correlations between connected nodes by the assortativity coefficient $r$, see Appendix A. This coefficient measures the tendency for nodes in a network to connect preferentially to others with similar degrees: When $r>0$ high-degree nodes tend to link to other high-degree nodes, while for $r<0$ high-degree nodes are more likely to connect with low-degree nodes. When $r=0$, degrees of connected nodes are independent, and thus the network is uncorrelated, following $P(k|k^{\prime})=k/\bar{k}P(k)$ Pastor-Satorras et al. (2015). Our approach is to perturbatively expand $P(k|k^{\prime})$ around the uncorrelated case at the first order in $r$, see Appendix A. We find that the effective reproduction number is approximated by
$R\approx R_{\mathrm{pert}}=(1-r)(R_{0}-m_{1})+r\left(\frac{\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}}{R_{0}-m_{1}}\right)\,,$ (6)
where $c_{k}$ is the probability that two neighbors of a node of degree $k$ are themselves connected, $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$ is the average number of triangles a link is part of, and $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$, see Appendix A.
Table 1: Reproduction numbers. Here, $R$ is the numerically computed leading eigenvalue of Eq. (3). The expressions of $R_{0}$ and $R_{\mathrm{pert}}$ are given by Eq. (5) and Eq. (6), respectively. To assess the impact of assortativity and clustering, we also compute $R_{\mathrm{pert}}$ for $r=0$ and for $\bar{c}=\sum_{k}c_{k}P(k)=0$. BAm: Barabási-Albert with parameter $m$. $\text{ER}_{d}$: Erdős-Rényi with mean degree $d$. WSk,p: Watts-Strogatz starting from a 1D lattice with $k$ nearest neighbors and rewiring probability $p$. Additional examples are given in sup .
@@
We compare $R_{\mathrm{pert}}$ given by Eq. (6) with the maximum eigenvalue $R$ of the matrix $M_{kk^{\prime}}$ for different network models, see Table 1. The relative difference between $R$ and $R_{\mathrm{pert}}$ is within $2\%$ for all the models we considered, except for Barabási-Albert with parameter $m=1$ ($20.6\%$) and $m=2$ ($7.1\%$), where $m$ represents the number of nodes a new node attaches to. This is likely due to the perturbative term being large in these two cases. Approximating $R$ with $R_{0}$ as given by Eq. (5) leads to substantially larger errors, supporting that clustering and assortativity play an important role. Here and in the following, when computing $R_{\mathrm{pert}}$ using Eq. (6) and $R_{0}$ using Eq. (5) for a given network, we always interpret the averages as empirical averages. With this choice, moments of the degree distribution remain finite even in very heterogeneous networks.
In practical cases, Eq. (6) predicts that $R_{\mathrm{pert}}$ increases with the assortativity $r$. Accordingly, assortativity lowers the epidemic threshold, as previously known for the unclustered case Goltsev et al. (2012). On the contrary, clustering substantially decreases $R_{\mathrm{pert}}$. Although assortativity and clustering influence the reproduction number in opposing ways, clustering typically exerts a stronger impact. This explains why $R_{0}$ consistently overestimates the effective reproduction number $R$ across all synthetic networks considered in our analysis. An exception occurs in networks with heavy-tailed degree distributions, where the influence of assortativity is significantly amplified, as $r$ approximately scales with $\overline{k^{3}}$ in Eq. (6), and may outweigh clustering effects (see Fig. 2 for an example).
Figure 2: Phase diagram representing the interplay between assortativity and clustering on the reproduction number for different degree distributions. We numerically compare the formulas for $R_{0}$, Eq. (5), and $R_{\rm pert}$, Eq. (6), with (a) a Poisson distribution with mean $5$ and (b) with a power-law with an exponential cutoff of $k_{c}=300$ ($P(k)\propto k^{-3}e^{-k/k_{c}}$). We assume that $c_{k}=m_{0}(k-1)^{-\alpha}$ and we compute $R_{\rm pert}$ using Eq. (6), for various values of $m_{0}$ and $\alpha$. The black curve represents the set of points characterized by $R_{\mathrm{pert}}=R_{0}$. The blue region represents the regime where $R_{\rm pert}$ is lower, whereas the red region represents the regime where $R_{\rm pert}$ increases the reproduction number compared to $R_{0}$.
IV.2 Case $p_{\psi}<1$
[…]
In our model, the infection spreads through every link at most once. A case in which the infection fails to spread via a specific link is equivalent to the case in which the link did not exist in the first place. This means that the scenario $p_{\psi}<1$ can be studied by removing links with uniform probability $(1-p_{\psi})$. This procedure alters the degree distribution of the network, which becomes
$P^{\psi}(k)=\sum_{q}\binom{q}{k}{p_{\psi}}^{k}\left(1-p_{\psi}\right)^{q-k}P(q).$ (7)
When the network is unassortative and unclustered, the reproduction number scales linearly with $p_{\psi}$:
$\sum_{k}(k^{2}-k)P^{\psi}(k)$ $={p_{\psi}}^{2}\left(\overline{k^{2}}-\overline{k}\right)$ (8)
$\sum_{k}kP^{\psi}(k)$ $={p_{\psi}}\overline{k}$ (9)
$\implies R^{\psi}=p_{\psi}R,$ (10)
where $R=\overline{k^{2}}/\overline{k}-1$, and the overbar still denotes averages over $P(k)$. Thus, the epidemic spreading condition ($R^{\psi}>1$) implies
$p_{\psi}>\frac{\overline{k}}{\overline{k^{2}}-\overline{k}}.$ (11)
This relation recovers the classic result for the epidemic threshold on uncorrelated networks Cohen et al. (2000); Newman (2002a).
This linear relation given in Eq. (8) does not hold, in general, for clustered or assortative networks. Triangles are vulnerable to link removal since all three links need to remain for a triangle to be preserved. If a given link belongs on average to $m_{1}$ triangles, after link removal it would belong on average to only $m_{1}{p_{\psi}}^{2}$ triangles. Therefore, the reproduction number, in the absence of degree correlations, is expressed by
$R^{\psi}=p_{\psi}R_{0}-m_{1}{p_{\psi}}^{2}.$ (12)
We now consider the effect of assortativity. Under edge removal with uniform probability $1-p_{\psi}$, the assortativity $r$ is transformed into
$r^{\psi}=r\left(1+\frac{p_{\psi}}{1-p_{\psi}}\frac{R_{0}}{\sigma_{r}^{2}}\right)^{-1},$ (13)
where $\sigma_{r}^{2}=\overline{k^{3}}/\overline{k}-\left(\overline{k^{2}}/\overline{k}\right)^{2}$. Equation (13) is derived in Ref. Noh (2007). Similarly to the uncorrelated case, we obtain
$\frac{\sum_{k}P^{\psi}(k)k(k-1)^{2}}{\sum_{k^{\prime}}k^{\prime}P^{\psi}(k^{\prime})}=p_{\psi}^{2}\frac{\overline{k(k-1)^{2}}}{\overline{k}}+R_{0}p_{\psi}(1-p_{\psi})$ (14)
and
$m_{2}^{\psi}\to p_{\psi}^{3}m_{2}+p_{\psi}^{2}(1-p_{\psi})m_{1}.$ (15)
Therefore, we have
$R_{\mathrm{pert}}^{\psi}$ $=p_{\psi}\left(R_{0}-m_{1}{p_{\psi}})\right)(1-r^{\psi})+r^{\psi}(1-p_{\psi})+$
$\frac{r^{\psi}}{R_{0}-m_{1}{p_{\psi}}}\left(p_{\psi}\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2p_{\psi}^{2}m_{2}\right.$
$\quad-p_{\psi}m_{1}(1-p_{\psi}+m_{1}p_{\psi}^{2})\biggr{)}+\mathcal{O}(r^{2})\,,$ (16)
where $r^{\psi}$ is given by Eq. (13).
V Network Euler-Lotka equation
We now use our estimates of $R$ to predict how epidemics spread in time on a network. The average number $\dot{I}_{k}(t)$ of newly infected individuals of degree $k$ at time $t$ is governed by the renewal equation